OpenStudy (anonymous):

If f(x) = 1 over 2x + 3 and g(x) = -4x - 1, find g[f(4)].

OpenStudy (anonymous):

\[f(4)=\frac{1}{2\times 4+3}=\frac{1}{11}\]

OpenStudy (anonymous):

it is -15/11

OpenStudy (anonymous):

g[f(4)] = -4* (1 over 2x +3) -1 = -4 * (1 over 2*4 +3) -1 = -4 * (1 over 11) -1 = -4/11 -1 = -15/11

OpenStudy (anonymous):

\[g(\frac{1}{11})=-4\times \frac{1}{11}-1\] what sheg said

OpenStudy (anonymous):

\[g(f(4))=g\left(\frac{1}{2\cdot4+3}\right)=g\left(\frac{1}{11}\right)=-4\cdot\frac{1}{11}-1=\frac{-4-11}{11}=-\frac{15}{11}\]

OpenStudy (anonymous):

But, one of the answer have to be. 21 -21 -19 19

OpenStudy (anonymous):

\[f(x) = \frac{1}{2x} + 3\] is it like this

OpenStudy (anonymous):

.... What?

OpenStudy (anonymous):

\[f(x)= 1/2 x+3 and g(x)= -4x-4 find g[f(4)]\]